# Learning production functions for supply chains with graph neural networks

### v1

@@ C.1 Production learning experiments
If product $p_{i}$ is a part of product $p_{o}$, then we would also expect that the products are close to each other in embedding space. To test this, we train node2vec embeddings (Grover and Leskovec,, 2016) on the firm-product graph, which is a static bipartite graph where an edge between a firm and a product indicates that the firm supplies or buys the product. We use embedding dimension 64 for the node2vec embeddings. Then, the score for product pair $p_{1}$, $p_{2}$ is the cosine similarity of their node2vec embeddings.
@@ C.2 Edge prediction experiments
As discussed in the main text, for all datasets, we order transactions chronologically and split them into train (70%), validation (15%), and test (15%). In these experiments, we train the models on the train set, choose hyperparameters based on the validation set, and report final results on the test set. When training the models, we train for a maximum of 100 epochs, but we allow for early stopping, if the validation MRR does not improve for over 10 epochs. In our main experiments (Table 2), we tested the following models: Edgebank, Static, Graph Transformer, SC-TGN, SC-TGN+inv, SC-GraphMixer, and SC-Graphmixer+inv. In Table 4, we report the hyperparameters we used for SC-TGN and SC-GraphMixer, which we found by sweeping over the hyperparameters. For all experiments, we ran 10 trials, each with a different random seed, and reported the mean and standard deviation in test MRR and test RMSE. As shown by the small standard deviations, performance is quite stable over random seeds, which we also show in Figures 7-8. These figures also show that SC-TGN and SC-Graphmixer both make steady progress in train and validation metrics over epochs: train loss decreases, validation MRR increases, and validation RMSE decreases. It is reassuring to see that the improvements in train loss correspond well to improvements in validation MRR and RMSE, since the train loss (16) combines multiple objectives, and, furthermore, we cannot directly train on MRR, which is non-differentiable, so instead we use softmax cross-entropy as a proxy loss.
@@
Table 5: Ablations of SC-TGN: comparing to original TGN (Rossi et al.,, 2020) and SC-TGN with ID embeddings, i.e., use memory directly as embedding, instead of applying GNN to memories. We only evaluate edge existence here, with mean reciprocal rank (MRR $\uparrow$), and leave out edge weight, since the original TGN did not predict edge weight. We report mean and standard deviation (in parentheses) over 10 seeds.
@@
We can view the Static and Graph Transformer baselines as ablations of our models, so we used the SC-TGN hyperparameters wherever applicable for fair comparison (e.g., embedding dimension, batch size). Since SC-TGN learns the initial memory, the Static baseline is equivalent to SC-TGN if no memory updates are applied and the embedding module is ID. The Static baseline is also a nested version of SC-GraphMixer, when we allow SC-GraphMixer to learn its initial node features. The Graph Transformer baseline is equivalent to SC-TGN if no memory updates are applied and the embedding module is UniMP (Shi et al.,, 2021). Since SC-TGN was our best-performing model, we also ran further ablations of SC-TGN to test the value of our extensions. We compared it to the original TGN, where we stripped away most of the extensions discussed in Appendix A.1: it no longer predicted edge weight, we removed the update penalty and learnable initial memory, and we trained it only on perturbation negatives. We also compared it to a version of SC-TGN that applied memory updates, but used ID for the embedding module, i.e., uses the node memory directly instead of applying a GNN to the memories. As shown in Table 5, we found that SC-TGN greatly outperformed both of these ablations as well.
@@
For the +inv experiments, we found that validation performance was best if we first pretrained the inventory module on only inventory loss (6), then co-trained the inventory module and SC-TGN / SC-GraphMixer, following the joint loss (16). Furthermore, in the +inv experiments, we experimented with whether to include the inventory module’s edge prediction penalties or not (8)-(9). Based on validation performance, we found that the penalties were useful on synthetic data, in the standard setting and under supply shocks, but not when there were missing transactions, or for the real-world datasets, so for those, we turned off the penalties. We believe this is because, while the inventory module is robust to missing transactions when learning production functions (as shown by our production learning results in Table 1), the penalties are not robust to missing transactions, since they rely on having an approximately correct and complete inventory. For example, if product $p_{i}$ is an input part to product $p_{o}$, and we are missing transactions where firm $f$ receives $p_{i}$, then the inventory module will incorrectly believe that $f$ has zero inventory of $p_{i}$ and penalize transactions where $f$ is supposed to supply $p_{o}$.

### v3

@@ C.1 Production learning experiments
If product $p_{i}$ is a part of product $p_{o}$, then we would also expect that the products are close to each other in embedding space. To test this, we train node2vec embeddings (Grover and Leskovec 2016) on the firm-product graph, which is a static bipartite graph where an edge between a firm and a product indicates that the firm supplies or buys the product. We use embedding dimension 64 for the node2vec embeddings. Then, the score for product pair $p_{1}$, $p_{2}$ is the cosine similarity of their node2vec embeddings.
@@ C.2 Edge prediction experiments
SS-std SS-shocks SS-missing Tesla IED
Edgebank (avg) 0.341 0.387 0.349 1.148 0.489
Static 0.343 (0.008) 0.425 (0.019) 0.374 (0.027) 1.011 (0.007) 0.504 (0.018)
SC-TGN 0.303 (0.003) 0.359 (0.007) 0.313 (0.002) 0.796 (0.012) 0.428 (0.011)
Table 5: Results for predicting weight of future edges (i.e., transaction amount), evaluated with root mean squared error (RMSE $\downarrow$). Standard scaling was applied to amounts, so an RMSE of 1.0 should be interpreted as being off by one standard deviation. We report mean and standard deviation (in parentheses) over 10 seeds.
@@
As discussed in the main text, for all datasets, we order transactions chronologically and split them into train (70%), validation (15%), and test (15%). In these experiments, we train the models on the train set, choose hyperparameters based on the validation set, and report final results on the test set. When training the models, we train for a maximum of 100 epochs, but we allow for early stopping, if the validation MRR does not improve for over 10 epochs. In our main experiments (Table 2), we tested the following models: Edgebank, Static, Graph Transformer, SC-TGN, SC-TGN+inv, SC-GraphMixer, and SC-Graphmixer+inv. In Table 6, we report the hyperparameters we used for SC-TGN and SC-GraphMixer, which we found by sweeping over the hyperparameters. We explore different model hyperparameters, such as dimensionality (e.g., memory and embedding dimensions for SC-TGN, node feature dimensions for SC-GraphMixer) and number of neighbors for node embedding, and different training hyperparameters, such as learning rate and batch size.
For all experiments, we ran 10 trials, each with a different random seed, and reported the mean and standard deviation in test MRR and test RMSE. As shown by the small standard deviations, performance is quite stable over random seeds, which we also show in Figures 6-7. These figures also show that SC-TGN and SC-Graphmixer both make steady progress in train and validation metrics over epochs: train loss decreases, validation MRR increases, and validation RMSE decreases. It is reassuring to see that the improvements in train loss correspond well to improvements in validation MRR and RMSE, since the train loss (16) combines multiple objectives, and, furthermore, we cannot directly train on MRR, which is non-differentiable, so instead we use softmax cross-entropy as a proxy loss.
@@
Table 7: Ablations of SC-TGN: comparing to original TGN (Rossi et al. 2020) and SC-TGN with ID embeddings, i.e., use memory directly as embedding, instead of applying GNN to memories. We only evaluate edge existence here, with mean reciprocal rank (MRR $\uparrow$), and leave out edge weight, since the original TGN did not predict edge weight. We report mean and standard deviation (in parentheses) over 10 seeds.
@@
We can view the Static and Graph Transformer baselines as ablations of our models, so we used the SC-TGN hyperparameters wherever applicable for fair comparison (e.g., embedding dimension, batch size). Since SC-TGN learns the initial memory, the Static baseline is equivalent to SC-TGN if no memory updates are applied and the embedding module is ID. The Static baseline is also a nested version of SC-GraphMixer, when we allow SC-GraphMixer to learn its initial node features. The Graph Transformer baseline is equivalent to SC-TGN if no memory updates are applied and the embedding module is UniMP (Shi et al. 2021). Since SC-TGN was our best-performing model, we also ran further ablations of SC-TGN to test the value of our extensions. We compared it to the original TGN, where we stripped away most of the extensions discussed in Appendix A.1: it no longer predicted edge weight, we removed the update penalty and learnable initial memory, and we trained it only on perturbation negatives. We also compared it to a version of SC-TGN that applied memory updates, but used ID for the embedding module, i.e., uses the node memory directly instead of applying a GNN to the memories. As shown in Table 7, we found that SC-TGN greatly outperformed both of these ablations as well.
@@
For the +inv experiments, we found that validation performance was best if we first pretrained the inventory module on only inventory loss (6), then co-trained the inventory module and SC-TGN / SC-GraphMixer, following the joint loss (16). Furthermore, in the +inv experiments, we experimented with whether to include the inventory module’s edge prediction penalties or not (14)-(15). Based on validation performance, we found that the penalties were useful on synthetic data, in the standard setting and under supply shocks, but not when there were missing transactions, or for the real-world datasets, so for those, we turned off the penalties. We believe this is because, while the inventory module is robust to missing transactions when learning production functions (as shown by our production learning results in Table 1), the penalties are not robust to missing transactions, since they rely on having an approximately correct and complete inventory. For example, if product $p_{i}$ is an input part to product $p_{o}$, and we are missing transactions where firm $f$ receives $p_{i}$, then the inventory module will incorrectly believe that $f$ has zero inventory of $p_{i}$ and penalize correct transactions where $f$ is supposed to supply $p_{o}$.